OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a).

OpenStudy (anonymous):

what is the actual question?

OpenStudy (anonymous):

oh make up your own?

OpenStudy (anonymous):

Yes lol

OpenStudy (anonymous):

then make one up!

OpenStudy (anonymous):

oh i see, you have to make one that can be "simplified" i.e. written in simplest radical form

OpenStudy (anonymous):

since \(5^2=25\) and \(2\times 25=50\) how about starting with \[\sqrt{50x^4}\]

OpenStudy (anonymous):

then you can "simplify" it as \[5x^2\sqrt{2}\]

OpenStudy (anonymous):

or make up one of your own

OpenStudy (anonymous):

Ok so, I would just put sqrt(50x^4) then say once it is simplified it becomes, 5x^2sqrt(2)

OpenStudy (anonymous):

you certainly could

OpenStudy (anonymous):

Thank you veryy much!